How much 10% acid solution and 30% acid solution should be mixed together to yield 100 liters of a 25% acid solution?
Answer by mananth(16946)

percent percentage ---------------- quantity
Acid I 10 ---------------- x liters
Acid II 30 ---------------- 100 - x liters
Mixture 25.00% ---------------- 100

10x+30(100-x)=25 * 100

10x +3000 -30x= 2500
10x-30x =2500 - 3000
-20 x= -500
/ -20
x=25 liters 10.00% Acid I
75 liters 30.00% Acid II